# Aligning LLMs to Ask Good Questions A Case Study in Clinical Reasoning

### v1

@@ Abstract
Large language models (LLMs) often fail to ask effective questions under uncertainty, making them unreliable in domains where proactive information-gathering is essential for decision-making. We present Alfa, a framework that improves LLM question-asking by (i) decomposing the notion of a “good” question into a set of theory-grounded attributes (e.g., clarity, relevance), (ii) controllably synthesizing attribute-specific question variations, and (iii) aligning models via preference-based optimization to explicitly learn to ask better questions along these fine-grained attributes. Focusing on clinical reasoning as a case study, we introduce the MediQ-AskDocs dataset, composed of 17k real-world clinical interactions augmented with 80k attribute-specific preference pairs of follow-up questions, as well as a novel expert-annotated interactive healthcare QA task to evaluate question-asking abilities. Models aligned with Alfa reduce diagnostic errors by 56.6% on MediQ-AskDocs compared to SOTA instruction-tuned LLMs, with a question-level win-rate of 64.4% and strong generalizability. Our findings suggest that explicitly guiding question-asking with structured, fine-grained attributes offers a scalable path to improve LLMs, especially in expert application domains. [^1] We release all data, code, and models for further research.
$\star$\$$\star$\$footnotetext: Equal contribution.
@@ 1 Introduction
Interactive language models have demonstrated remarkable capabilities across numerous domains (OpenAI et al., 2024), yet proactive interaction abilities in high-stakes scenarios—clinical reasoning, legal analysis, investigative journalism—remains a challenge (Fung et al., 2024). A key obstacle is the ability of these models to recognize and anticipate missing or ambiguous information and proactively seek clarification (Li et al., 2024; Deng et al., 2024). In clinical practice, for instance, physicians systematically ask patients questions to rule out or confirm relevant diagnoses (Richardson et al., 1995; Proffit, 2013). This iterative, information-seeking behavior is essential for accurate and safe decision-making. Similarly, for large language models (LLMs) to serve as reliable decision-support tools for clinicians, they must learn not only to provide answers, but also to identify when additional information is needed, and to ask follow-up questions that effectively reduce uncertainty (Figure 1).
However, there are two main challenges in building LLMs that ask good questions, especially in expert domains. First, defining a “good” question is inherently complex and context-dependent. In general, attributes such as clarity, focus, and answerability are essential (Heritage and Maynard, 2006; Roter and Hall, 1987; Freed, 1994; Searle, 1969); however, in domain-specific scenarios such as clinical reasoning, additional properties—medical accuracy, diagnostic relevance, and mitigating differential diagnosis (DDX) biases—are necessary for reducing diagnostic uncertainty (Richardson et al., 1995; Silverman et al., 2016; Heritage, 2010; Hall et al., 1995; West, 1984; Stivers and Majid, 2007; Ong et al., 1995). Second, instilling the ability to ask good questions in LLMs is technically non-trivial (Li et al., 2024; Johri et al., 2025; Zhang et al., 2024a). Naïve prompting strategies such as “Ask a follow-up question if needed.” may enhance interactivity but lack a principled foundation for defining a good question. We propose leveraging well-established general and task-specific principles from communication theory and psychology to improve LLMs’ information-seeking abilities.
@@
    Decompose the high-level goal of asking “good” questions into structured, grounded attributes.
@@
    Synthesize counterfactual data by controllably altering specific attributes (e.g. make clearer).
@@
We instantiate the above recipe with a focus on clinical reasoning, where question-asking is central to reducing diagnostic uncertainty and preventing errors. To this end, we construct a novel dataset, MediQ-AskDocs, containing 17k clinical interactions with follow-up questions from the r/AskDocs subreddit, paired with 80k synthesized counterfactual variants of these questions highlighting each attribute. These counterfactual pairs provide fine-grained training signals for preference learning. Finally, we integrate the attribute-specific signals into a unified policy by combining all synthetic data into a single model, training separate reward models and merging them, or fusing attribute-specific policies. Alfa contrasts with coarse-grained preference learning, offering a more targeted way to refine question-asking behaviors.
@@ 2 Problem Statement
Aligning LLMs to ask good question requires contextual reasoning along multiple attributes (e.g., accuracy, clarity, focus). However, most alignment paradigms treat these goals as monolithic, aggregating preferences into a single reward that conflates attributes and obscures their individual contributions. We formalize this challenge as follows:
@@
$\mathcal{D}_{synth}^{k}=\{(a_{i}^{k+},a_{i}^{k-})|R_{k}(a_{i}^{k+})>R_{k}(a_{i}^{k-})\}.\vspace{-1mm}$
@@
$\vspace{-1mm}\pi^{*}=\arg\max_{\pi}\mathbb{E}_{(s,a)\sim\pi}[f(R_{1},\ldots,R_{K})(s,a)],\vspace{-2mm}$
@@ 3 What Makes a “Good” Question?
To systematically improve LLM question asking, we define six key attributes grounded in cognitive science, psychology, and clinical communication research (Heritage and Maynard, 2006; Roter and Hall, 1987; Chouinard, 2007; Freed, 1994; Searle, 1969; Levinson, 2012). Unlike prior work that relies on implicit heuristics, we explicitly decompose question quality into interpretable, tangible attributes that enhance clinical reasoning.
General Question-Quality Attributes.
Effective questions must be clear, targeted, and answerable to drive meaningful interactions. We select three core attributes:
  • Clarity: Avoids ambiguity and unnecessary complexity, ensuring precise communication (Heritage and Maynard, 2006; Roter and Hall, 1987).
  • Focus: Directly addresses a specific information gap, yielding more informative responses (Ronfard et al., 2018; Gopnik and Wellman, 2012; Chouinard, 2007; Freed, 1994). E.g., “Do you have a family history of breast cancer?” is superior to “Has anyone in your family been sick?”
  • Answerability: Adapts to the respondent’s expertise (e.g., no jargans) (Levinson, 2012; Keil et al., 2008; Searle, 1969; Zhu and Wu, 2025).
Domain-specific question-asking attributes.
In clinical reasoning, question-asking is a structured diagnostic skill. Drawing from clinical communication research (Richardson et al., 1995; Silverman et al., 2016; Heritage, 2010; Hall et al., 1995; West, 1984; Stivers and Majid, 2007; Ong et al., 1995; Proffit, 2013), we define three additional attributes essential for clinical decision-making:
  • Medical Accuracy: Aligns with established medical knowledge & guidelines.
  • Diagnostic Relevance: Probes for symptoms, risk factors, or contextual details essential to refining differential diagnoses (DDX).
  • Avoiding DDX Bias: Prevents suggestive or leading wording that could introduce cognitive biases and misguide diagnostic reasoning.
@@ 4.2 Attribute-Specific Data Generation
Counterfactual Perturbation.
For each question $a_{i}$ in the dataset, we prompt an LLM to create “enhanced” and “corrupted” variants $(a_{i}^{k+},a_{i}^{k-})$ that explicitly alter only one attribute $k$ at a time (e.g., rewriting the original question to be more clear/ambiguous) while keeping others consistent. This enables us to create controlled preference pairs, where one version of a question $a_{i}^{k+}$ is better aligned with a specific attribute than the other, $a_{i}^{k-}$.
Verification & Filtering.
We use an LLM-judge to verify that the generated perturbations reflect their intended modification. With original question $a_{i}$, we obtain an enhanced question $a_{i}^{k+}$ and a corrupted question $a_{i}^{k-}$, resulting in three preference pairs: $(a_{i}^{k+},a_{i})$, $(a_{i}^{k+},a_{i}^{k-})$, and $(a_{i},a_{i}^{k-})$. Given each pair, we provide additional context [^2] Parsed conversation conclusions from future turns. to the LLM-judge, and ask the judge to compare the pairs in the specified attribute dimension (e.g., which question is clearer). If the judge’s decision matches the intended perturbation direction—verifying that $R_{k}(a_{i}^{k+})>R_{k}(a_{i}^{k-})$—we retain the sample; otherwise, we discard it. This filtering step removes inconsistencies and ensures that our synthetic preference data provides reliable supervision for model alignment. See details in Appendix C.1.
@@ 4.3 Attribute Integration Strategies
We now combine signals from the attribute-specific data so that the model produces questions that optimize for all attributes. Standard preference optimization algorithms DPO and PPO allow distinct Points of Integration (POI) as described below [^3] While reward sum, which trains attribute-specific reward models then combines the reward scores (e.g., via a learned linear combination or average) to produce a final scalar reward (Wu et al., 2023; Wang et al., 2024c), remains an option, we do not utilize this strategy due to high compute cost of loading all reward models during PPO without substantial performance improvements (Rame et al., 2024; Shi et al., 2024).:
(1) Data Mixing pools all synthetic data into one training set and uses standard DPO/PPO. This treats each attribute-specific comparison as part of a larger set of “better vs. worse” question pairs, enabling a single policy to learn from all attributes at once (Wang et al., 2023; Lambert et al., 2024)
(2) Reward Fusion trains separate reward models, then averages the reward model weights in an overall reward model which can then be used in PPO to align a final policy (Ramé et al., 2024).
(3) Policy Fusion trains separate policies or preference models, each specialized for one attribute, and then combine the model weights by averaging or taking a linear combination (Jang et al., 2023). This strategy offers the highest degree of parallelism and maximally preserves attribute strengths and interpretability.
@@ 5.2 Experiments
With the goal of instilling question-asking ability in LLMs in the clinical reasoning domain, we structure our experiments to progressively address two key questions: (1) Does the entire Alfa pipeline improve clinical question-asking performance? and (2) Necessity of each component of Alfa and to what extent do they contribute to the performance? Accordingly, we organize our evaluation into three parts:
@@
  • We compare the decomposed attributes (clarity, medical accuracy, etc.) to a “coarse” attribute (simply “good or bad”) to highlight the necessity of a theory-grounded decomposition (§6.2).
@@
3. Ablation Studies. We dissect each design choice to identify its role in the observed improvements. Specifically, we compare attribute-integration strategies6.4), apply quality filter on the synthetic data (§6.5), examine data perturbation directions (corruption vs. enhancement) (§6.6), ablate individual attributes6.7), and test out-of-distribution generalization6.8) on a separate medical reasoning task. [^4] All ablations are done with DPO due to its lighter compute requirements unless otherwise specified.
@@ 5.3 Evaluation
We create an LLM-judge to compare pairs of questions, adopting prompt structures from Li et al. (2023) and Dubois et al. (2023). Specifically, we measure the percentage of times our aligned models’ questions are preferred over those of the baseline instruction-tuned models (llama-3.2-3b-Instruct and llama-3.1-8b-Instruct) by to the judge and report win-rate. Each comparison is carried out by gpt-4o with rationales, permuting the order of the questions in each pair to mitigate ordering bias. As an sanity check, we find that the LLM-judge assigns a higher win-rate to questions crafted by verified experts than to those written by non-experts, suggesting that our evaluation aligns with domain expertise.
Expert manual evaluation.   To assess the quality of generated questions and further validate the LLM-judge, we collect preference rankings from medical experts and compute win-rate of select models. See Appendix F for further details.
Accuracy on interactive diagnostic task.
@@
MediQ-AskDocs Task.   MediQ is compatible with any QA task with contextual information. We introduce a novel healthcare QA task as part of the MediQ-AskDocs dataset: 302 consumer healthcare multiple choice questions manually annotated by medical experts. The task is automatically generated by o1 using the test split of MediQ-AskDocs5.1), and annotated by two human experts with 88.6% agreement. See Appendix E for task construction and annotation details. Additionally, we use MedQA (Jin et al., 2020) with MediQ to examine the models’ ability to generalize to out-of-domain tasks.
@@ 6 Results & Analysis
The results section first demonstrates the superior performance of Alfa6.1). Then we investigate how the core modules—attribute decomposition6.2) and preference tuning6.3)—contributes to its effectiveness. Additionally, we ablate on attribute integration strategies6.4), synthetic data quality6.5), direction of synthetic data (corruption vs. enhancement) (§6.6), attribute-specific influences6.7), and the out-of-distribution generalization6.8) of Alfa.
@@ 6.1 Alfa Improves Overall Question Asking
We begin by comparing Alfa with two baselines: (1) the base models (llama-3.2-3B-Instruct and llama-3.1-8B-Instruct) and (2) models trained via supervised fine-tuning (SFT) on the human-written questions. The Policy Fusion attribute integration strategy (§4.3) is reported for both Alfa-DPO and Alfa-PPO in Table 1 below.
@@
Table 1: Main results. Alfa models consistently outperform base instruct models and SFT models.
We first confirm that fine-tuning the human-written questions in MediQ-AskDocs (SFT) already outperforms base models, establishing usefulness of the dataset. Notably, performance of the Alfa-aligned models shows explicitly modeling structured, theory-grounded attributes substantially boosts both question quality and diagnostic accuracy, reducing diagnostic errors by 56.62%.
6.2 Fine-Grained Attributes Outperform Coarse Attribute
The core idea of Alfa is to decompose a complex goal into structured, theory-grounded attributes rather than treating it as one coarse objective. We examine the role of attribute decomposition by comparing alignmend with fine-grained attributes (Alfa) with a simpler approach that optimizes models on a coarse “better” vs. “worse” distinction. Both models undergo the same counterfactual data generation and alignment processes, but the coarse model lacks explicit attribute separation.
@@ 6.2 Fine-Grained Attributes Outperform Coarse Attribute
Table 2: Fine-grained (Alfa) vs. Coarse Attributes. Fine-grained attributes lead to better downstream diagnostic accuracy and similar win-rates compared to coarse-grained objective. Expert evaluation for 3B coarse vs. fine-grained shows identical win-rate: 59.4%.
Table 2 shows that Alfa, trained with fine-grained attributes, achieves higher performance compared to models trained on coarse attributes on MediQ diagnostic accuracy, while the two methods show comparable LLM-judge win-rates and identical expert evaluation win-rates. Aligning models based on structured attributes guides the reasoning process in a theory-grounded way, resulting in substantially better downstream performance. Additionally, qualitative analysis (Figure 3) shows that the coarse-aligned model tends to ask more superficial questions such as waiting for 8 weeks and check back, while the Alfa-model’s question, "Was this the only issue that was found?", indicates some form of reasoning to rule out other factors that might have contributed to the patient’s concern. Further, as we show in §6.8, coarsely aligned models generalize poorly to out-of-distribution tasks.
Figure 3: Quanlitative example contrasting models aligned with fine-grained and coarse attributes. Alfa model tends to ask more logical questions.

### v2

# Alfa: Aligning LLMs to Ask Good Questions A Case Study in Clinical Reasoning

@@ Abstract
Large language models (LLMs) often fail to ask effective questions under uncertainty, making them unreliable in domains where proactive information-gathering is essential for decision-making. We present ALignment via Fine-grained Attributes, (Alfa) a framework that improves LLM question-asking by (i) decomposing the notion of a “good” question into a set of theory-grounded attributes (e.g., clarity, relevance), (ii) controllably synthesizing attribute-specific question variations, and (iii) aligning models via preference-based optimization to explicitly learn to ask better questions along these fine-grained attributes. Focusing on clinical reasoning as a case study, we introduce the MediQ-AskDocs dataset, composed of 17k real-world clinical interactions augmented with 80k attribute-specific preference pairs of follow-up questions, as well as a novel expert-annotated interactive healthcare QA task to evaluate question-asking abilities. Models aligned with Alfa reduce diagnostic errors by 56.6% on MediQ-AskDocs compared to SoTA instruction-tuned LLMs, with a question-level win-rate of 64.4% and strong generalizability. Our findings suggest that explicitly guiding question-asking with structured, fine-grained attributes offers a scalable path to improve LLMs, especially in expert application domains. [^1] We release all data, code, and models for further research.
@@ 1 Introduction
Interactive language models have demonstrated remarkable capabilities across numerous domains (OpenAI et al., 2024), yet proactive interaction abilities in high-stakes scenarios—clinical reasoning, legal analysis, investigative journalism—remains a challenge (Fung et al., 2024). A key obstacle is the ability of these models to recognize and anticipate missing or ambiguous information and proactively seek clarification (Li et al., 2024; Deng et al., 2024). In clinical practice, for instance, physicians systematically ask patients questions to rule out or confirm relevant diagnoses (Richardson et al., 1995; Proffit, 2013). This iterative, information-seeking behavior is essential for accurate and safe decision-making. Similarly, for large language models (LLMs) to serve as reliable decision-support tools for clinicians, they must learn not only to provide answers, but also to identify when additional information is needed, and to ask follow-up questions that effectively explore and adjust possible hypothesis or reduce uncertainty (Figure 1).
However, there are two main challenges in building LLMs that ask good questions, especially in expert domains. First, defining a “good” question is inherently complex and context-dependent. In general, attributes such as clarity, focus, and answerability are essential (Heritage & Maynard, 2006; Roter & Hall, 1987; Freed, 1994; Searle, 1969); however, in domain-specific scenarios such as clinical reasoning, additional properties—medical accuracy, diagnostic relevance, and mitigating differential diagnosis (DDX) biases—are necessary for reducing diagnostic uncertainty (Richardson et al., 1995; Silverman et al., 2016; Heritage, 2010; Hall et al., 1995; West, 1984; Stivers & Majid, 2007; Ong et al., 1995). Second, instilling the ability to ask good questions in LLMs is technically non-trivial (Li et al., 2024; Johri et al., 2025; Zhang et al., 2024a). Naïve prompting strategies such as “Ask a follow-up question if needed.” may enhance interactivity but lack a principled foundation for defining a good question. We propose leveraging well-established general and task-specific principles from communication theory and psychology to improve LLMs’ information-seeking abilities.
@@
    Decompose the goal of asking “good” questions into structured, grounded attributes.
@@
    Synthesize counterfactual data by controllably altering any attribute (e.g. make clearer).
@@
We instantiate the above recipe with a focus on clinical reasoning, where question-asking is central to reducing diagnostic uncertainty and preventing errors. To this end, we construct a novel dataset, MediQ-AskDocs, containing 17k clinical interactions with follow-up questions from the r/AskDocs subreddit [^2] With medical experts verified following subreddit policy. See §Limitations for further discussion on setting., paired with 80k synthesized counterfactual variants of these questions highlighting each attribute. These counterfactual pairs provide fine-grained training signals for preference learning (Li et al., 2025). Finally, we integrate the attribute-specific signals into a unified policy by combining all synthetic data into a single model, training separate reward models and merging them, or fusing attribute-specific policies. Alfa contrasts with coarse-grained preference learning, offering a more targeted way to refine question-asking behaviors.
@@ 2 Problem Statement
Aligning LLMs to ask good question requires reasoning along multiple attributes (e.g., accuracy, clarity, focus). However, most alignment paradigms treat these goals as monolithic, aggregating preferences into a single reward that conflates attributes and obscures their individual contributions. We formalize this challenge as follows:
@@
$\mathcal{D}_{synth}^{k}=\{(a_{i}^{k+},a_{i}^{k-})|R_{k}(a_{i}^{k+})>R_{k}(a_{i}^{k-})\}.\vskip-2.84526pt$
@@
$\vskip-2.84526pt\pi^{*}=\arg\max_{\pi}\mathbb{E}_{(s,a)\sim\pi}[f(R_{1},\ldots,R_{K})(s,a)],\vskip-2.84526pt$
@@ 3 What Makes a “Good” Question?
To systematically improve LLM question asking, we define six key attributes grounded in cognitive science, psychology, and clinical communication research (Heritage & Maynard, 2006; Roter & Hall, 1987; Chouinard, 2007; Freed, 1994; Searle, 1969; Levinson, 2012). Unlike prior work that relies on implicit heuristics, we explicitly decompose question quality into interpretable, tangible attributes that enhance clinical reasoning.
General Question-Quality Attributes. Effective questions must be clear, targeted, and answerable to drive meaningful interactions. We select three core attributes:
  1. 1.
    Clarity, aiming to avoid ambiguity and unnecessary complexity (e.g., no jargons), ensuring precise communication (Heritage & Maynard, 2006; Roter & Hall, 1987; Burns et al., 2022);
  2. 2.
    Focus, directly addressing a specific information gap, yielding more informative responses (Ronfard et al., 2018; Gopnik & Wellman, 2012; Chouinard, 2007; Freed, 1994). E.g., “Has anyone in your family had breast cancer?” is superior to “Has anyone in your family been sick?”; and
  3. 3.
    Answerability, ensuring the question is both within the respondent’s knowledge domain and appropriate for them to answer (e.g., asking about their symptoms and experiences rather than expecting them to provide medical diagnoses or knowledge that falls within the clinician’s expertise) (Levinson, 2012; Keil et al., 2008; Searle, 1969).
Domain-specific question-asking attributes. In clinical reasoning, question-asking is a structured diagnostic skill. Drawing from clinical communication research (Richardson et al., 1995; Silverman et al., 2016; Heritage, 2010; Hall et al., 1995; West, 1984; Stivers & Majid, 2007; Ong et al., 1995; Proffit, 2013), we define three additional attributes essential for clinical decision-making:
  1. 4.
    Medical Accuracy requires alignment with established medical textbook knowledge & guidelines;
  2. 5.
    Diagnostic Relevance probes for symptoms, risk factors, or contextual details essential to refining differential diagnoses (DDX); and
  3. 6.
    Avoiding DDX Bias prevents suggestive or leading wording that could introduce cognitive biases and misguide diagnostic reasoning.
@@ 4.2 Attribute-Specific Data Generation
Counterfactual Perturbation. For each question $a_{i}$ in the dataset, we prompt an LLM to create “enhanced” and “corrupted” variants $(a_{i}^{k+},a_{i}^{k-})$ that explicitly alter only one attribute $k$ at a time (e.g., rewriting the original question to be more clear/ambiguous) while keeping others consistent. This enables us to create controlled preference pairs, where one version of a question $a_{i}^{k+}$ is better aligned with a specific attribute than the other, $a_{i}^{k-}$.
Verification & Filtering. We use an LLM-judge (Zheng et al., 2023) to verify that the generated perturbations reflect their intended modification. With original question $a_{i}$, we obtain an enhanced question $a_{i}^{k+}$ and a corrupted question $a_{i}^{k-}$, resulting in three preference pairs: $(a_{i}^{k+},a_{i})$, $(a_{i}^{k+},a_{i}^{k-})$, and $(a_{i},a_{i}^{k-})$. Given each pair, we provide additional context [^3] Parsed conversation conclusions from future turns. to the LLM-judge, and ask the judge to compare the pairs in the specified attribute dimension (e.g., which question is clearer). If the judge’s decision matches the intended perturbation direction—verifying that $R_{k}(a_{i}^{k+})>R_{k}(a_{i}^{k-})$—we retain the sample; otherwise, we discard it. This filtering step removes inconsistencies and ensures that our synthetic preference data provides reliable supervision for model alignment. See details in Appendix C.1.
@@ 4.3 Attribute Integration Strategies
We now combine signals from the attribute-specific data so that the model produces questions that optimize for all attributes. Standard preference optimization algorithms DPO (Rafailov et al., 2024) and PPO (Schulman et al., 2017) allow distinct Points of Integration (POI) as described below [^4] While reward sum, which trains attribute-specific reward models then combines the reward scores (e.g., via a learned linear combination or average) to produce a final scalar reward (Wu et al., 2023; Wang et al., 2024c), remains an option, we do not utilize this strategy due to high compute cost of loading all reward models during PPO without substantial performance improvements (Rame et al., 2024; Shi et al., 2024).:
  1. (1)
    Data Mixing pools all data into one training set and uses standard DPO/PPO. This treats each attribute-specific comparison as part of a larger set of “better vs. worse” pairs, enabling a single policy to learn from all attributes at once (Wang et al., 2023b; Lambert et al., 2024).
  2. (2)
    Reward Fusion trains separate reward models (RM), then averages the RM weights in an overall RM which can then be used in PPO to align a final policy (Ramé et al., 2024).
  3. (3)
    Policy Fusion trains separate policies or preference models, each specialized for one attribute, and then combine the model weights by averaging or taking a linear combination (Jang et al., 2023). This strategy offers the highest degree of parallelism and maximally preserves attribute strengths and interpretability.
@@ 5.2 Experiments
With the goal of instilling question-asking ability in LLMs in the clinical reasoning domain, we structure our experiments to progressively address two key questions: (1) Does the entire Alfa pipeline improve clinical question-asking performance? and (2) To what extent is every component of Alfa necessary for improving performance? Accordingly, we organize our evaluation as follows:
@@
  • We compare the decomposed attributes (e.g., clarity, answerability) to a “coarse” attribute (simply “good or bad”) to examine the effect of theory-grounded decomposition (§6.2).
@@
3. Ablation Studies. We dissect each design choice to identify its role in the observed improvements. We compare attribute-integration strategies, apply quality filter on the synthetic data, examine data perturbation directions (corruption vs. enhancement), ablate individual attributes, and test out-of-distribution generalization on a separate diagnostic task [^5] All ablations are done with DPO due to lighter compute requirements unless otherwise specified.6.4).
@@ 5.3 Evaluation
We create an LLM-judge to compare pairs of questions, adopting prompt structures from Li et al. (2023) and Dubois et al. (2023). Specifically, we measure the percentage of times our aligned models’ questions are preferred over those of the baseline instruction-tuned models (llama-3.2-3b-Instruct and llama-3.1-8b-Instruct) by the judge and report win-rate. Each comparison is carried out by gpt-4o with rationales, permuting the presentation order in each pair to mitigate ordering bias. As an validity check, the LLM-judge assigns a higher win-rate to questions from verified experts than those from non-experts (Appendix D), suggesting our evaluation aligns with domain expertise.
Expert manual evaluation.
To assess the quality of generated questions and further validate the LLM-judge, we conduct manual preference rankings with three medical experts from our research team and compute win-rate of select models based on majority vote. See Appendix F for further details.
Interactive diagnostic accuracy.
@@
The MediQ-AskDocs task.
MediQ is compatible with any QA task with contextual information. We introduce a novel healthcare QA task as part of the MediQ-AskDocs dataset: 302 consumer healthcare multiple choice questions manually annotated by medical experts. The task is automatically generated by o1 using the test split of MediQ-AskDocs5.1) and achieves 85.9% agreement with majority voted manual annotations from medical experts. See Appendix E for task construction and annotation details. Additionally, we use MedQA (Jin et al., 2020) with MediQ to examine the models’ ability to generalize out-of-domain.
@@ 6.1 Alfa Improves Overall Question Asking
We begin by comparing Alfa with two baselines: (1) the base instruction-tuned models (llama-3.2-3B-Instruct and llama-3.1-8B-Instruct) and (2) models trained via supervised fine-tuning (SFT) on the human-written questions. The Policy Fusion attribute integration strategy (§4.3) is reported for both Alfa-DPO and Alfa-PPO in Table 1.
Our results confirm that fine-tuning the human-written questions in MediQ-AskDocs (SFT) already outperforms base models, establishing usefulness of the dataset. Notably, superior performance of the Alfa-aligned models shows explicitly modeling structured, theory-grounded attributes substantially boosts both question quality and diagnostic accuracy, reducing diagnostic errors by 56.62% (increasing accuracy by 21.5%).
@@
Table 1: Main results. Alfa models consistently outperform base instruct and SFT models. Note that win-rates for base models are set to 50% to represent equal preference when comparing against themselves.
@@
Table 2: Fine-grained (Alfa) vs. Coarse Attributes. Fine-grained attributes lead to better downstream diagnostic accuracy and similar win-rates compared to coarse-grained objective. Expert evaluation for 3B coarse vs. fine-grained shows identical win-rate against the base model: 59.4%.
6.2 Fine-Grained Attributes Outperform Coarse Attribute
The core idea of Alfa is to decompose a complex goal into structured, theory-grounded attributes rather than treating it as one coarse objective. We examine the role of attribute decomposition by comparing alignment with fine-grained attributes (Alfa) with a simpler approach that optimizes models on a coarse “better” vs. “worse” distinction (Geng et al., 2025). Both models undergo the same counterfactual data generation and alignment processes, but the coarse model lacks explicit attribute separation.
Figure 3: Qualitative example contrasting models aligned with fine-grained and coarse attributes. Alfa model tends to ask more logical questions.
Results in Table 2 show that Alfa, trained with fine-grained attributes, achieves higher performance compared to models trained on coarse attributes on MediQ diagnostic accuracy, while the two methods show comparable LLM-judge win-rates. Aligning models based on structured attributes guides the reasoning process in a theory-grounded way, resulting in substantially better downstream performance. Additionally, qualitative analysis (Figure 3) shows that the coarse-aligned model tends to ask more superficial questions such as waiting for 8 weeks and check back, while the Alfa-model’s question, "Was this the only issue that was found?", indicates some form of reasoning to rule out other factors that might have contributed to the patient’s concern. Further, coarsely aligned models generalize poorly to out-of-distribution tasks (§6.4).